OpenStudy (anonymous):

A quiz has 5 true-false questions. How many ways can a student answer all the questions?

OpenStudy (destinymasha):

2 choices , 5 questions. 2 x 5 =10

OpenStudy (anonymous):

Thanks! I thought that was it, but I wasn't sure.

jimthompson5910 (jim_thompson5910):

More like, there are 2*2*2*2*2 = 2^5 = 32 different ways to answer all the questions.

OpenStudy (anonymous):

So which is it???????

OpenStudy (destinymasha):

maybe i was wrong....

OpenStudy (anonymous):

I'll give the answer in a second. Fingers crossed, it's right!!

OpenStudy (destinymasha):

which one?

OpenStudy (anonymous):

It was 32!

OpenStudy (destinymasha):

oops...

OpenStudy (anonymous):

It's okay!